H purchases components from three suppliers. Components purchased from A are priced at $4.5 each and used at the rate of 20,000 units per year. Components purchased from B are priced at $5 each and are used at the rate of 3,500 units per year. Components purchased from C are priced at $4.7 each and used at the rate of 1,000 units per year. Currently, H purchases a separate truckload from each supplier. As part of its JIT drive, H has decided to aggregate purchases from the three suppliers. The trucking company charges a fixed cost of $550 for the truck with an additional charge of $80 for each stop. Thus, if H asks for a pickup from only one supplier, the trucking company charges $630; from two suppliers, it charges $710; and from three suppliers, it charges $790. Suggest a replenishment strategy (lots ordered independently or jointly?) for Harley that minimizes annual cost. Assume a holding cost of 25% per year.
Answer #1

A replenishment strategy where lots are ordered jointly minimizes annual cost.

Component A Component B ComponentC
Demand, Da 20000 Demand, Db 3500 Demand, Dc 1000
Shipping pickup, S = Fixed cost (550) + one stop (80) $             630.00 Shipping pickup, S $         630.00 Shipping pickup, S $       630.00
Price $                  4.50 Price $             5.00 Price $            4.70
Holding cost, Ha $                1.125 Holding cost, Hb $             1.00 Holding cost, Hc $            0.94
(25% of the price) (25% of the price) (25% of the price)
EOQ = Sqrt(2DS/H) 4732.86 EOQ 2100.00 EOQ 1157.77
Current Strategy - Ordered individually
Cycle inventory = D/EOQ 4.23 1.67 0.86
Inventory cost = (EOQ*H)/2 $          2,662.24 Inventory cost = (EOQ*H)/2 $     1,050.00 Inventory cost = (EOQ*H)/2 $       544.15
Order cost = (D/EOQ) * 500 $          2,662.24 Order cost = (D/EOQ) * 500 $     1,050.00 Order cost = (D/EOQ) * 500 $       544.15
Total inventory cost current strategy= 2662.24+1050+544.15 $       4,256.390
Total order cost current strategy $       4,256.390
Total cost Current strategy $          8,512.78 (4256.39+4256.39)
New Shipping cost, Sn = 790 $             790.00
Optimal joint order frequency, N
n* = SQRT((DAHa + DBHb + DcHC)/(2S*))
DAHa + DBHb + DcHC = 26940
2S              1,580.00
n*= sqrt (26940/1580) 4.13
Order Quantity for A, Qa = Da/n* 4843.50 Order Quantity for B, Qb = Db/n* 847.61 Order Quantity for C, Qc = Dc/n* 242.18
Inventory cost = (Q*H)/2 $          2,724.47 Inventory cost = (Q*H)/2 $         423.81 Inventory cost = (Q*H)/2 $       113.82
Total inventory cost new strategy $          3,262.10
Total order cost new strategy= 790*4.13 $          3,262.10
Saving in Inventory cost $         994.29
Saving in ordering cost $         994.29
Total Savings $     1,988.58